find the present value of a 2 year deffered annuity at 4% interest compounded quarterly with payments of P1100.00 made every quarter for 3 years


1
Expert's answer
2021-02-03T02:18:04-0500

We can use the below present value formula

"PV = pmt x\frac{ [(1 - \frac{1} {(1 + i)}^n)]}{ i}"

PV = present value = $?

Pmt = Periodic payment = $1100

i = interest rate = 4% = 0.04

n = Number of remaining payments = 2 years

so n = 2 x 4 = 8 payments

we can use below formula

"PV = pmt x\frac{ [(1 - \frac{1} {(1 + i)}^n)]}{ i}"

"PV = 1100\frac{ [(1 - \frac{1} {(1 + 0.004)}^8)]}{ 0.004}"

PV = $8646

So the present value is $8646